OpenStudy (anonymous):

√3x+6 (+6 is in radical)= x-4

OpenStudy (anonymous):

square both sides to get rid of square root

OpenStudy (anonymous):

square (x-4)---FOIL

OpenStudy (anonymous):

\[3x + 6 = x^{2} -8x + 16\]

OpenStudy (anonymous):

subtract 3x and 6 to move to right side leaving 0 on the left

OpenStudy (anonymous):

\[x^{2} -11x - 10 =0\]

OpenStudy (anonymous):

you can factor using chart or my grouping

OpenStudy (anonymous):

sorry that should be +10

OpenStudy (anonymous):

\[x^{2} - 11x + 10 =0\]

OpenStudy (anonymous):

what factors of 10 as a product gives you -11 as a sum? -1 * -10=10 and -1 + (-10)=-11

OpenStudy (anonymous):

so the factors are -1 and -10 (x-1)(x-10)=0 x-1 = 0 and x-10=0 x=1 x=10 {1, 10}

OpenStudy (anonymous):

or factoring by grouping-break down middle term using know factors of -1 and-10

OpenStudy (anonymous):

\[x^{2}-x-10x+10\]

OpenStudy (anonymous):

factor out GCF x(x-1)-10(x-1)=0 (x-10)(x-1)=0 x=10 and x=1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome...do you understand? what grade are you in...medal please ; )

OpenStudy (anonymous):

im in 8th doing algebra, so what ur final answer?

OpenStudy (anonymous):

and what are you talking about with a medal?

OpenStudy (anonymous):

love 8th grade Algebra! Solutions are 10 and 1 {1, 10}

OpenStudy (anonymous):

click on good answer by my name on the right

OpenStudy (anonymous):

yeah i got that, but when you plug it back in, it doesnt work. does it have to work?

OpenStudy (anonymous):

what answer do i click on for good answer?

OpenStudy (anonymous):

it does work

OpenStudy (anonymous):

\[x=1 \sqrt{9}=1-4\]

OpenStudy (anonymous):

okay good, thanks for your help

OpenStudy (anonymous):

square root of 9 is 3 or -3

OpenStudy (anonymous):

wait could u also help me with another question instead of me reposting one??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x=10 so square root of 36 equals -6 or 6

OpenStudy (anonymous):

a right triangle has one leg that is 4 inches shorter than the other leg. the hypotenuse is √106 inches. find the lengths of the legs.

OpenStudy (anonymous):

ok i'm writing it out to solve...hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ill be right back

OpenStudy (anonymous):

do you want to factor my grouping

OpenStudy (anonymous):

\[2x^{2}-8x-90=0\]

OpenStudy (anonymous):

let a = x, b=(x-4) and c=square root of 106(hypotenuse)

OpenStudy (anonymous):

\[x^{2} + x^{2} -8x +16 = 106\]

OpenStudy (anonymous):

after plugging into Pythagorean Theorem (right triangle)

OpenStudy (anonymous):

subtract 106 from both sides to get the first equation I posted above

OpenStudy (anonymous):

back

OpenStudy (anonymous):

working backwards so hope you can follow

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

where did you get lost?

OpenStudy (anonymous):

http://www.twiddla.com/536745 go there for explanation

OpenStudy (anonymous):

thx